OpenStudy (anonymous):

Anyone know how to do this one Find the x- and y- intercepts. If there are no intercepts write "No x-intercept" or "No y-intercept". y=2/3x-16

OpenStudy (anonymous):

at the x-intercept, y is 0, so substitute 0 in for y : (0)=2/3x-16 16=2/3x x=24 Now do the same to find the y-intercept, but this time substitute 0 in for x to find y.

OpenStudy (anonymous):

kk thax